Question 1 2 pts What is the present value of $100 ten years from now if the simple rate of interest is0.25? Round your final answer to two decimal places. Question 2 2 pts A deposit of $11 is made at the beginning of each year for 17 years, what is the future value of all deposits if simple rate of interest is 0.11? Round your final answer to two decimal places. Question 3 2 pts In how many years will $100 become $320 if simple rate of interest is 11%? 20
Answer #1

Question 1

Amount = $ 100

Simple interest rate = 0.25

Time = 10 years

Present value = Amount / ((1+r)^n)

=$ 100 / ((1.25) ^ 10)

= $ 10.74

Question 2

Future value Amountrate time Amount Rate Years Future value 3157 3036 2915 2794 2673 2552 2431 231 2189 2068 1947 1826 1705 15 84 1463 1342 1221 37213 11 01117 11 01116 11 011 15 11 01114 11 011 13 11 011 12 11 01110 11 011 11 011 11 011 11 011 11 011 11 011 11 011 11 011 11 011 9 7 4 TOTAL FUTURE VALUE

Question 3

Present value = $ 100

Maturity value = $ 320

Time = ?

Rate of interest = 11%

Formula used = Interest = Maturity value – Present value

And Interest = Present value * Time * Rate

Therefore, Interest = 320-100 = $ 220

Time = Interest / ( present value * Rate of interest)

Time = 220 / ( 100*0.11)

Time = 20 years